# Policy Learning with Abstention

### v2

@@ 1 Introduction
While abstaining algorithms have been developed in the classification literature (bousquet2021fast), generalizing these algorithms to the setting of policy learning from observational data is a non-trivial task. The primary difficulty is the counterfactual nature of the problem: the learner only ever observes an outcome under treatment or control for any given unit, never both. Because of this, the learner must transform samples (either via inverse propensity weighting or through a doubly-robust correction) into “pseudo-outcomes” — quantities representing the unit-specific contrasts in outcomes between treatment and control.
@@
    Also in Section 3, we introduce our primary algorithm for policy learning with abstention (Algorithm 1). Our algorithm first de-biases observations via inverse propensity weighting (IPW) when propensities are known (Section 3.1), or by performing a doubly-robust correction when propensities are unknown (Section 3.2). We prove the algorithm obtains fast $O(1/n)$ high probability regret rates under the safe/abstaining regret, and in the worst-case obtains $O(1/\sqrt{n})$ standard regret, matching rates due to athey2021policy.
@@
In addition to the above, we also conduct simulations to compare our safe policy improvement algorithm (Algorithm 3) to relevant benchmarks (thomas2015high). In sum, our work illustrates the practicality and importance of studying policy learning with abstention, and further provides insight into how abstention can impact other areas of policy learning as well.
@@
Classification with abstention/a reject option traces back to chow1970optimum, who characterize the optimal error–reject trade-off, prescribing abstention whenever the posterior risk of mis-classification exceeds the reject cost. herbei2006reject consider classification with a reject option in a statistical learning setting, and bartlett2008reject introduce a convex “reject-hinge” loss to make abstention trainable in large-scale settings. Other researchers have developed abstaining classification algorithms for SVMs (grandvalet2008svmreject), multiclass classification (ramaswamy2015consistent), boosting (cortes2016boosting), online learning  (cortes2018online; pasteris2024bandits; neu2020fast; vanderHoeven2022regretvariance), and even deep learning settings (geifman2017selective; geifman2019selectivenet). madras2018defer study a setting where the prediction of an abstaining algorithm is replaced by a downstream expert, but don’t provide regret/excess risk guarantees. Most closely related to this paper is the work of bousquet2021fast, who develop an empirical risk minimization-based algorithm for abstention that enables fast $O(1/n)$ risk minimization rates without standard margin/realizability assumptions.
@@
Our work builds off of the literature on policy learning, particularly works on welfare maximization/regret minimization (athey2021policy; hirano2009asymptotics; kitagawa2018should; manski2004statistical). Also related to our work is the literature of safe policy improvement, which aims to produce a policy that, with high-probability, improves over some baseline “safe” policy. Existing approaches include using hypothesis testing to select between the baseline and learned policy (thomas2015high; cho2025cspi), bootstrap the baseline policy with a learned policy in regions of low confidence (laroche2019safe; simao2019safe), and directly minimizing “negative regret” (ghavamzadeh2016safe). Our work provides new perspective onto safe policy improvement, showing how abstaining algorithms can naturally be used to construct improved policies. Other complementary areas that are not directly considered in the work are policy evaluation (dudik2011doubly; karampatziakis2021off), offline reinforcement learning (moodie2007demystifying; moodie2012q; murphy2003optimal), policy learning under partial identification (kallus2018confounding), and inference on values of optimal treatment policies (luedtke2016statistical; chen2023inference; whitehouse2025inference).
@@ 2 Notation and Preliminaries
A policy $\pi:{\mathcal{X}}\to\{0,1\}$ prescribes a binary treatment to a unit with covariates $X$. Let $\Pi$ denote a class of feasible treatment policies. We assume the $\Pi$ has bounded complexity, controlled through the VC dimension (vapnik2013nature).
@@ 3 Policy Learning with Abstention
$g_{o}(*,x)\coloneqq\mathbb{E}\Big[\tfrac{Y(1)+Y(0)}{2}+p\,\Big|\,X=x\Big].$ (2)
@@
$V^{(p)}(\pi)$ $\coloneqq\mathbb{E}\!\Big[\mathbbm{1}\!\!\left\{\pi(X)\neq*\right\}\,v(\pi,X)\;+\;\mathbbm{1}\!\!\left\{\pi(X)=*\right\}\,g_{o}(*,X)\Big].$ (3)
@@
+ $\qquad+\;\mathbbm{1}\!\!\left\{\pi(X)=*\right\}\big(\tfrac{YD}{2p_{o}(X)}+\tfrac{Y(1-D)}{2(1-p_{o}(X))}+p\big)\Big].$ (4)
@@ 3.1 Known Propensities
Theorem 3.1 analyzes policy learning with an abstention bonus and establishes an $O(1/n)$ fast rate when performance is measured against the best binary policy in the class. This benchmark is standard in the abstention literature (see, e.g., (bousquet2021fast, Section 1.2)). In Section 4.3, we will discuss a natural connection between this regret bound and distributional robustness.
@@
When the additive bonus is $p=0$, our algorithm can no longer be expected to obtain fast $O(1/n)$ rates. The following proposition shows that, in this undiscounted setting, Algorithm 1 still obtains $O(1/\sqrt{n})$ regret rates outlined, which are known to be generally unimprovable without margin or realizability assumption (athey2021policy). In this setting, we can convert an abstaining policy into a binary one by having it assign a treatment uniformly at random in the set $\{0,1\}$. The following proposition provides this slower regret rate, showing there is no added risk of using our algorithm over just empirical welfare maximization even in setting without the additive bonus. We prove this result in Appendix A.
@@ 3.2 Unknown Propensities: Doubly Robust Learner
Taking inspiration from the literature on semiparametric estimation (chernozhukov2018double; bang2005doubly; foster2023orthogonal; athey2021policy), we introduce a doubly-robust analogue of Algorithm 1 that furnishes fast regret rates even when the propensity is unknown. Our algorithm uses ML estimates $\widehat{g},\widehat{p}$ of the regression $g_{o}(d,x)$ and propensity $p_{o}(x)$ to “de-bias” observed outcomes $Y$ into more robust pseudo-outcomes. These nuisance estimates are assumed to be independent of the sample. In more detail, we define
@@ 4.1 Fast Learning Rates Without Standard Margin Assumptions
In policy learning, fast regret rates are sometimes achievable when the CATE is deterministically bounded away from zero, as outlined in the margin condition of Equation (6). In particular, prior works have shown that $O(n^{-1})$ rates are possible in the realizable setting, where the Bayes-optimal policy $\pi^{B}$ lies in the class $\Pi$ (kitagawa2018should; Luedtke2020). However, it is generally impossible to know if a policy class $\Pi$ (say, all depth 3 decision trees) contains the optimal treatment policy in advance. Thus, one should aim to obtain fast regret rates under in agnostic settings, where $\Pi$ may not contain the optimal treatment strategy.
In the classification literature, recent works achieve fast learning rates in agnostic settings (ben2014sample; bousquet2021fast). These works eschew the realizability assumption, instead imposing a finite combinatorial diameter on the policy class. In words, a class of policies $\Pi$ which bounds the maximal number of points of disagreement between any two policies in the class. More formally, the combinatorial diameter is defined by
@@
Theorem 4.2 unifies the regression-rate guarantee (Luedtke2020) and fast rate in the realizable case (kitagawa2018should). In particular, when $\beta=1/2$, one recovers $n^{-1}$ regret. More generally, if $\beta<1/2$ but the policy class $\Pi$ is expressive enough that $V(\pi^{B}(X))-V(\pi^{*}(X))$ is small (e.g., $\leq c/n^{-\alpha}$ for some $\alpha>0$), the resulting bound improves upon a direct plug-in classifier ($\pi(X)=\mathbbm{1}\!\!\left\{\hat{\tau}(X)>0\right\}$). Thus, either a strong CATE oracle or an “almost” realizable policy class is sufficient to surpass the $1/\sqrt{n}$ regret.
@@ 4.2 Safe Policy Improvement
Safe policy improvement (SPI) addresses the deployment problem under uncertainty: one should adopt a new policy only if it can be certified to outperform a baseline policy $\omega$ with high probability. Following recent work on safe policy improvement thomas2015high; cho2025cspi, we use sample splitting (into $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{test}}$) to separate the tasks of candidate policy selection and safety testing. This splitting enables one-sided lower confidence bounds (LCBs) on the improvement over the baseline: $V(\hat{\pi})-V(\omega)$. We treat Algorithm 1 as a black box to propose a $\{0,1,*\}$-valued policy $\widetilde{\pi}$} learned on $\mathcal{D}_{train}$. To obtain a deployable binary policy, we impute abstentions with the baseline policy to obtain a candidate policy Equation (7).
@@
We compare Algorithm 3 with an Empirical Welfare Maximizer (EWM) and prior work on synthetic data. In particular, we consider the two variants of High-Confidence Policy Improvement (HCPI) described in thomas2015high: one based on finite-sample confidence intervals and one based on a $t$-test statistic. For EWM, we choose the empirical welfare maximizer as the candidate policy and test it against the baseline in exactly the same way as in Algorithm 3 (with $k=1$).
@@ Appendix A Missing Proofs from Section˜3
First, we restate that a standard result regarding the VC subgraph dimension (see Lemma A.1 of kitagawa2018should):
@@ A.1 Concentration Inequalities
The inequality Equation˜9 holds trivially for any $g$ satisfying $\mathbb{E}g^{2}<r(n,\delta)^{2}$. Now, if $\mathbb{E}g^{2}\geq r(n,\delta)^{2}$, define
@@
This gives us the bound in Equation˜9.
@@
To bound the covering numbers, we use the fact that the covering numbers of a star-shaped hull $\mathcal{G}$ of a class $\mathcal{F}$ (of VC dimension $d$) satisfy (bartlett2006empirical):
@@
The second-to-last inequality follows from standard bounds on the covering numbers of VC classes. Substituting $\gamma=d/n$ into Equation˜11, we get
@@
Taking expectation in Equation˜12 we have
@@
The inequality above, together with Equation˜10, implies that
@@
Finally, by applying Talagrand’s inequality (Theorem 3.27 of wainwright2019high), we obtain
@@
Now we invoke the Peeling Argument (wainwright2019high). For $k=0,1,\dots,K:=\lceil\log_{2}(n)\rceil$, define shells $\mathcal{G}_{k}=\{g\in\mathcal{G}:2^{-(k+1)}<\mathbb{E}g^{2}\leq 2^{-k}\}$. Apply the tail bound to each $\mathcal{G}_{k}$ with confidence levels $\delta_{k}=\delta/(K+1)$ and union bound over $k$. On the resulting $1-\delta$ event we have, simultaneously for all $g\in\mathcal{G}$,
@@
has VC dimension is at most $10d$ bousquet2021fast; vidyasagar2013learning. We now invoke Lemma˜A.2 with $\Pi=\mathcal{H}$ , $\gamma_{1}(z)=\left|\frac{\kappa\ y\,d}{p_{o}(x)}+\frac{\kappa y\,(1-d)}{1-p_{o}(x)}\right|$, $\gamma_{2}(z)=0$ and $B=2$. ∎
@@
Substituting $p=0$ in Equation˜14 and noting that ${\mathcal{D}}({\widehat{\Pi}})\leq\textbf{c}\,$ for some constant c , proves the result. ∎
@@ Appendix B Doubly Robust Abstention: Additional Details and Proofs
so the near–optimal set in Step 5 uses $\alpha_{\mathrm{DR}}$ and the disagreement between with the EWM $\mathbb{E}_{n}|{\widehat{\pi}}-\pi|$. Where $\mathrm{Err}_{\mathrm{DR}}$ is any upper bound on product error in the nuisance: $\mathrm{Err}_{\mathrm{DR}}\;\geq\;\mathbb{E}\!\Big[(\widehat{p}(X)-p_{o}(X))^{2}\textstyle\sum_{d=0}^{1}(\widehat{g}(d,X)-g_{o}(d,X))^{2}\Big]^{1/2}$. In practice, we simply increase the constant in Step 4 of Algorithm 1; when the nuisance product error is $o_{p}(n^{-1/2})$, this yields the same rate as Theorem 3.1. Such product rates are a common assumption in the doubly robust literature (foster2023orthogonal) and hold for several function classes as discussed in (foster2023orthogonal, Section 5 and Appendix E). Moreover, a natural choice of $\mathrm{Err}_{\mathrm{DR}}$ is the $L_{4}$ error rates:
@@
The minimax $L_{4}$ error rates are well-understood for many nonparametric classes of interest, such as smooth classes (stone1980optimal; stone1982optimal), Holder classes (lepskii1992asymptotically), Besov classes (donoho1998minimax) and convex function classes (guntuboyina2015global).

### v3

@@ 1 Introduction
While abstaining algorithms have been developed in the classification literature (Bousquet and Zhivotovskiy, 2021), generalizing these algorithms to the setting of policy learning from observational data is a non-trivial task. The primary difficulty is the counterfactual nature of the problem: the learner only ever observes an outcome under treatment or control for any given unit, never both. Because of this, the learner must transform samples (either via inverse propensity weighting or through a doubly-robust correction) into “pseudo-outcomes” — quantities representing the unit-specific contrasts in outcomes between treatment and control.
@@
    Also in Section 3, we introduce our primary algorithm for policy learning with abstention (Algorithm 1). Our algorithm first de-biases observations via inverse propensity weighting (IPW) when propensities are known (Section 3.1), or by performing a doubly-robust correction when propensities are unknown (Section 3.2). We prove the algorithm obtains fast $O(1/n)$ high probability regret rates under the safe/abstaining regret, and in the worst-case obtains $O(1/\sqrt{n})$ standard regret, matching rates due to Athey and Wager (2021).
@@
In addition to the above, we also conduct simulations to compare our safe policy improvement algorithm (Algorithm 3) to relevant benchmarks (Thomas et al., 2015). In sum, our work illustrates the practicality and importance of studying policy learning with abstention, and further provides insight into how abstention can impact other areas of policy learning as well.
@@
Classification with abstention/a reject option traces back to Chow (1970), who characterize the optimal error–reject trade-off, prescribing abstention whenever the posterior risk of mis-classification exceeds the reject cost. Herbei and Wegkamp (2006) consider classification with a reject option in a statistical learning setting, and Bartlett and Wegkamp (2008) introduce a convex “reject-hinge” loss to make abstention trainable in large-scale settings. Other researchers have developed abstaining classification algorithms for SVMs (Grandvalet et al., 2009), multiclass classification (Ramaswamy et al., 2015), boosting (Cortes et al., 2016), online learning  (Cortes et al., 2018; Pasteris et al., 2024; Neu and Zhivotovskiy, 2020; van der Hoeven et al., 2022), and even deep learning settings (Geifman and El-Yaniv, 2017, 2019). Madras et al. (2018) study a setting where the prediction of an abstaining algorithm is replaced by a downstream expert, but don’t provide regret/excess risk guarantees. Most closely related to this paper is the work of Bousquet and Zhivotovskiy (2021), who develop an empirical risk minimization-based algorithm for abstention that enables fast $O(1/n)$ risk minimization rates without standard margin/realizability assumptions.
@@
Our work builds off of the literature on policy learning, particularly works on welfare maximization/regret minimization (Athey and Wager, 2021; Hirano and Porter, 2009; Kitagawa and Tetenov, 2018; Manski, 2004). Also related to our work is the literature of safe policy improvement, which aims to produce a policy that, with high-probability, improves over some baseline “safe” policy. Existing approaches include using hypothesis testing to select between the baseline and learned policy (Thomas et al., 2015; Cho et al., 2025), bootstrap the baseline policy with a learned policy in regions of low confidence (Laroche et al., 2019; Simão et al., 2019), and directly minimizing “negative regret” (Ghavamzadeh et al., 2016). Our work provides new perspective onto safe policy improvement, showing how abstaining algorithms can naturally be used to construct improved policies. Other complementary areas that are not directly considered in the work are policy evaluation (Dudík et al., 2011; Karampatziakis et al., 2021), offline reinforcement learning (Moodie et al., 2007, 2012; Murphy, 2003), policy learning under partial identification (Kallus and Zhou, 2018), and inference on values of optimal treatment policies (Luedtke and Van Der Laan, 2016; Chen et al., 2023; Whitehouse et al., 2025).
@@ 2 Notation and Preliminaries
A policy $\pi:{\mathcal{X}}\to\{0,1\}$ prescribes a binary treatment to a unit with covariates $X$. Let $\Pi$ denote a class of feasible treatment policies. We assume the $\Pi$ has bounded complexity, controlled through the VC dimension (Vapnik, 2013).
@@ 3 Policy Learning with Abstention
$g_{o}(*,x)\coloneqq\mathbb{E}\Big[\tfrac{Y(1)+Y(0)}{2}+p\,\Big|\,X=x\Big].$
@@
$V^{(p)}(\pi)$ $\coloneqq\mathbb{E}\!\Big[\mathbbm{1}\!\!\left\{\pi(X)\neq*\right\}\,v(\pi,X)\;+\;\mathbbm{1}\!\!\left\{\pi(X)=*\right\}\,g_{o}(*,X)\Big].$
@@
+ $\qquad+\;\mathbbm{1}\!\!\left\{\pi(X)=*\right\}\big(\tfrac{YD}{2p_{o}(X)}+\tfrac{Y(1-D)}{2(1-p_{o}(X))}+p\big)\Big].$
@@ 3.1 Known Propensities
Theorem 3.1 analyzes policy learning with an abstention bonus and establishes an $O(1/n)$ fast rate when performance is measured against the best binary policy in the class. This benchmark is standard in the abstention literature (see, e.g., (Bousquet and Zhivotovskiy, 2021, Section 1.2)). In Section 4.3, we will discuss a natural connection between this regret bound and distributional robustness.
@@
When the additive bonus is $p=0$, our algorithm can no longer be expected to obtain fast $O(1/n)$ rates. The following proposition shows that, in this undiscounted setting, Algorithm 1 still obtains $O(1/\sqrt{n})$ regret rates outlined, which are known to be generally unimprovable without margin or realizability assumption (Athey and Wager, 2021). In this setting, we can convert an abstaining policy into a binary one by having it assign a treatment uniformly at random in the set $\{0,1\}$. The following proposition provides this slower regret rate, showing there is no added risk of using our algorithm over just empirical welfare maximization even in setting without the additive bonus. We prove this result in Appendix A.
@@ 3.2 Unknown Propensities: Doubly Robust Learner
Taking inspiration from the literature on semiparametric estimation (Chernozhukov et al., 2018; Bang and Robins, 2005; Foster and Syrgkanis, 2023; Athey and Wager, 2021), we introduce a doubly-robust analogue of Algorithm 1 that furnishes fast regret rates even when the propensity is unknown. Our algorithm uses ML estimates $\widehat{g},\widehat{p}$ of the regression $g_{o}(d,x)$ and propensity $p_{o}(x)$ to “de-bias” observed outcomes $Y$ into more robust pseudo-outcomes. These nuisance estimates are assumed to be independent of the sample. In more detail, we define
@@ 4.1 Fast Learning Rates Without Standard Margin Assumptions
In policy learning, fast regret rates are sometimes achievable when the CATE is deterministically bounded away from zero, as outlined in the margin condition of Equation (3). In particular, prior works have shown that $O(n^{-1})$ rates are possible in the realizable setting, where the Bayes-optimal policy $\pi^{B}$ lies in the class $\Pi$ (Kitagawa and Tetenov, 2018; Luedtke and Chambaz, 2020). However, it is generally impossible to know if a policy class $\Pi$ (say, all depth 3 decision trees) contains the optimal treatment policy in advance. Thus, one should aim to obtain fast regret rates under in agnostic settings, where $\Pi$ may not contain the optimal treatment strategy.
In the classification literature, recent works achieve fast learning rates in agnostic settings (Ben-David and Urner, 2014; Bousquet and Zhivotovskiy, 2021). These works eschew the realizability assumption, instead imposing a finite combinatorial diameter on the policy class. In words, a class of policies $\Pi$ which bounds the maximal number of points of disagreement between any two policies in the class. More formally, the combinatorial diameter is defined by
@@
Theorem 4.2 unifies the regression-rate guarantee (Luedtke and Chambaz, 2020) and fast rate in the realizable case (Kitagawa and Tetenov, 2018). In particular, when $\beta=1/2$, one recovers $n^{-1}$ regret. More generally, if $\beta<1/2$ but the policy class $\Pi$ is expressive enough that $V(\pi^{B}(X))-V(\pi^{*}(X))$ is small (e.g., $\leq c/n^{-\alpha}$ for some $\alpha>0$), the resulting bound improves upon a direct plug-in classifier ($\pi(X)=\mathbbm{1}\!\!\left\{\hat{\tau}(X)>0\right\}$). Thus, either a strong CATE oracle or an “almost” realizable policy class is sufficient to surpass the $1/\sqrt{n}$ regret.
@@ 4.2 Safe Policy Improvement
Safe policy improvement (SPI) addresses the deployment problem under uncertainty: one should adopt a new policy only if it can be certified to outperform a baseline policy $\omega$ with high probability. Following recent work on safe policy improvement Thomas et al. (2015); Cho et al. (2025), we use sample splitting (into $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{test}}$) to separate the tasks of candidate policy selection and safety testing. This splitting enables one-sided lower confidence bounds (LCBs) on the improvement over the baseline: $V(\hat{\pi})-V(\omega)$. We treat Algorithm 1 as a black box to propose a $\{0,1,*\}$-valued policy $\widetilde{\pi}$} learned on $\mathcal{D}_{train}$. To obtain a deployable binary policy, we impute abstentions with the baseline policy to obtain a candidate policy Equation (4).
@@
We compare Algorithm 3 with an Empirical Welfare Maximizer (EWM) and prior work on synthetic data. In particular, we consider the two variants of High-Confidence Policy Improvement (HCPI) described in Thomas et al. (2015): one based on finite-sample confidence intervals and one based on a $t$-test statistic. For EWM, we choose the empirical welfare maximizer as the candidate policy and test it against the baseline in exactly the same way as in Algorithm 3 (with $k=1$).
@@ Appendix A Missing Proofs from Section˜3
First, we restate that a standard result regarding the VC subgraph dimension (see Lemma A.1 of Kitagawa and Tetenov (2018)):
@@ A.1 Concentration Inequalities
The inequality Equation˜6 holds trivially for any $g$ satisfying $\mathbb{E}g^{2}<r(n,\delta)^{2}$. Now, if $\mathbb{E}g^{2}\geq r(n,\delta)^{2}$, define
@@
This gives us the bound in Equation˜6.
@@
To bound the covering numbers, we use the fact that the covering numbers of a star-shaped hull $\mathcal{G}$ of a class $\mathcal{F}$ (of VC dimension $d$) satisfy (Bartlett and Mendelson, 2006):
@@
The second-to-last inequality follows from standard bounds on the covering numbers of VC classes. Substituting $\gamma=d/n$ into Equation˜8, we get
@@
Taking expectation in Equation˜9 we have
@@
The inequality above, together with Equation˜7, implies that
@@
Finally, by applying Talagrand’s inequality (Theorem 3.27 of Wainwright (2019)), we obtain
@@
Now we invoke the Peeling Argument (Wainwright, 2019). For $k=0,1,\dots,K:=\lceil\log_{2}(n)\rceil$, define shells $\mathcal{G}_{k}=\{g\in\mathcal{G}:2^{-(k+1)}<\mathbb{E}g^{2}\leq 2^{-k}\}$. Apply the tail bound to each $\mathcal{G}_{k}$ with confidence levels $\delta_{k}=\delta/(K+1)$ and union bound over $k$. On the resulting $1-\delta$ event we have, simultaneously for all $g\in\mathcal{G}$,
@@
has VC dimension is at most $10d$ Bousquet and Zhivotovskiy (2021); Vidyasagar (2013). We now invoke Lemma˜A.2 with $\Pi=\mathcal{H}$ , $\gamma_{1}(z)=\left|\frac{\kappa\ y\,d}{p_{o}(x)}+\frac{\kappa y\,(1-d)}{1-p_{o}(x)}\right|$, $\gamma_{2}(z)=0$ and $B=2$. ∎
@@
Substituting $p=0$ in Equation˜11 and noting that ${\mathcal{D}}({\widehat{\Pi}})\leq\textbf{c}\,$ for some constant c , proves the result. ∎
@@ Appendix B Doubly Robust Abstention: Additional Details and Proofs
so the near–optimal set in Step 5 uses $\alpha_{\mathrm{DR}}$ and the disagreement between with the EWM $\mathbb{E}_{n}|{\widehat{\pi}}-\pi|$. Where $\mathrm{Err}_{\mathrm{DR}}$ is any upper bound on product error in the nuisance: $\mathrm{Err}_{\mathrm{DR}}\;\geq\;\mathbb{E}\!\Big[(\widehat{p}(X)-p_{o}(X))^{2}\textstyle\sum_{d=0}^{1}(\widehat{g}(d,X)-g_{o}(d,X))^{2}\Big]^{1/2}$. In practice, we simply increase the constant in Step 4 of Algorithm 1; when the nuisance product error is $o_{p}(n^{-1/2})$, this yields the same rate as Theorem 3.1. Such product rates are a common assumption in the doubly robust literature (Foster and Syrgkanis, 2023) and hold for several function classes as discussed in (Foster and Syrgkanis, 2023, Section 5 and Appendix E). Moreover, a natural choice of $\mathrm{Err}_{\mathrm{DR}}$ is the $L_{4}$ error rates:
@@
The minimax $L_{4}$ error rates are well-understood for many nonparametric classes of interest, such as smooth classes (Stone, 1980, 1982), Holder classes (Lepskii, 1992), Besov classes (Donoho and Johnstone, 1998) and convex function classes (Guntuboyina and Sen, 2015).